OpenStudy (anonymous):

I need the answer for 8(c-3)=96

Parth (parthkohli):

Start by dividing both sides by 8. \( \color{Black}{\Rightarrow c - 3 = 12 }\) Can you do it now?

OpenStudy (anonymous):

8c-24=96 8c=96+24 (move the -24 over to the other side which makes it a positive.) 8c=120 ( divide the 120 by 8) c=15

OpenStudy (anonymous):

TY I love the simple breakdown

OpenStudy (anonymous):

np :)